OpenStudy (anonymous):

Estimate to the nearest integer square root of 197

OpenStudy (jonnyvonny):

Integer, from what I remember, is whole number. So, the way I'd approach this is just go up with with x^2: 1*1=1 2*2=4... 12*12=144 13*13=169 14*14=196 15*15=225 The closet one is 13*13. So that is your integer; 13.